# Efficient 3D Instance Mapping and Localization with Neural Fields https://gtangg12.github.io/iML

### v2

@@ Abstract
We tackle the problem of learning an implicit scene representation for 3D instance segmentation from a sequence of posed RGB images. Towards this, we introduce 3DIML, a novel framework that efficiently learns a label field that may be rendered from novel viewpoints to produce view-consistent instance segmentation masks. 3DIML significantly improves upon training and inference runtimes of existing implicit scene representation based methods. Opposed to prior art that optimizes a neural field in a self-supervised manner, requiring complicated training procedures and loss function design, 3DIML leverages a two-phase process. The first phase, InstanceMap, takes as input 2D segmentation masks of the image sequence generated by a frontend instance segmentation model, and associates corresponding masks across images to 3D labels. These almost view-consistent pseudolabel masks are then used in the second phase, InstanceLift, to supervise the training of a neural label field, which interpolates regions missed by InstanceMap and resolves ambiguities. Additionally, we introduce InstanceLoc, which enables near realtime localization of instance masks given a trained label field and an off-the-shelf image segmentation model by fusing outputs from both. We evaluate 3DIML on sequences from the Replica and ScanNet datasets and demonstrate 3DIML’s effectiveness under mild assumptions for the image sequences. We achieve a large practical speedup over existing implicit scene representation methods with comparable quality, showcasing its potential to facilitate faster and more effective 3D scene understanding.
@@ I Introduction
  • An overall practical runtime improvement of 14-24$\times$ over prior art benchmarked on a single GPU (NVIDIA RTX 3090)
@@ II Background
2D segmentation: The prevalence of vision transformer architecture and the increasing scale of image datasets have resulted in a series of state-of-the-art image segmentation models. Panoptic and Contrastive Lifting both lift panoptic segmentation masks produced by Mask2Former [1] to 3D by learning a neural field. Towards open-set segmentation, segment anything (SAM) [2] achieves unprecedented performance by training on a billion masks over 11 million images. HQ-SAM [9] improves upon SAM for fine-grained masks. FastSAM [8] distills SAM into a CNN architecture and achieves similar performance while being orders of magnitude faster. In this work, we use GroundedSAM [10, 11], which refines SAM to produce object-level, as opposed to part-level segmentation masks.
Neural fields for 3D instance segmentation: NeRFs are implicit scene representations that can accurately encode complex geometry, semantics, and other modalities, as well as resolve viewpoint inconsistent supervision [12]. Panoptic lifting [5] constructs semantics and instances branches on an efficient variant of NeRF, TensoRF [13], utilizing a Hungarian matching loss function to assign learned instance masks to surrogate object IDs given reference view-inconsistent masks. This scales poorly with increasing number of objects (owing to the cubic complexity of Hungarian matching). Contrastive lifting [6] addresses this by instead employing contrastive learning on scene features, with positive and negative relations determined by whether or not they project onto the same mask. In addition, contrastive lifting requires a slow-fast clustering-based loss for stable training, leading to faster performance than panoptic lifting but requires multiple stages of training, leading to slow convergence. Concurrently to us, Instance-NeRF [14] directly learn a label field, but they based their mask association on utilizing NeRF-RPN [15] to detect objects in a NeRF. Our approach, on the contrary, allows scaling to very high image resolutions while requiring only a small number (40-60) of neural field queries to render segmentation masks.
Structure from Motion: During mask association in InstanceMap, we take inspiration from scalable 3D reconstruction pipelines such as hLoc [16], including the use of visual descriptors for matching image viewpoints first, then applying keypoint matching as a preliminary for mask association. We utilize LoFTR [17] for keypoint extraction and matching.
@@ IV-C Implementation Details
For comparison with Panoptic and Contrastive Lifting, we follow their evaluation framework and map the Mask2Former outputs and ground truth semantic labels to 21 reduced ScanNet classes, considering only the "thing" subset of those classes. We train Panoptic and Contrastive Lifting using the code provided by [6]. We use the default parameterizations as their Replica experiment, which scales the number of iterations (128000) with the number of images. We build 3DIML on top of the popular framework Nerfstudio [18] with tinycudann enabled [19]. We train until convergence for 5000 iterations for Replica-vMap (both Mask2Former and GroundedSAM) and 10000 iterations for ScanNet. For our panoptic experiments, we use $\left(\tau_{\text{global}},\tau_{\text{local}},\tau_{\text{community}},\tau_{\text{area}},\tau_{\text{merge}}\right)=$ (0.25, 0.8, 2, 0.15,0.75). For GroundedSAM, we use (0.25, 0.75, 3, 0.3, 0.6) for Replica-vMap and (0.4, 0.75, 3, 0.25, 0.75) for ScanNet.

### v5

@@ Abstract
We tackle the problem of learning an implicit scene representation for 3D instance segmentation from a sequence of posed RGB images. Towards this, we introduce 3DIML, a novel framework that efficiently learns a neural label field which can render 3D instance segmentation masks from novel viewpoints. Opposed to prior art that optimizes a neural field in a self-supervised manner, requiring complicated training procedures and loss function design, 3DIML leverages a two-phase process. The first phase, InstanceMap, takes as input 2D segmentation masks of the image sequence generated by a frontend instance segmentation model, and associates corresponding masks across images to 3D labels. These almost 3D-consistent pseudolabel masks are then used in the second phase, InstanceLift, to supervise the training of a neural label field, which interpolates regions missed by InstanceMap and resolves ambiguities. Additionally, we introduce InstanceLoc, which enables near realtime localization of instance masks given a trained neural label field. We evaluate 3DIML on sequences from the Replica and ScanNet datasets and demonstrate its effectiveness under mild assumptions for the image sequences. We achieve a large practical speedup over existing implicit scene representation methods with comparable quality, showcasing its potential to facilitate faster and more effective 3D scene understanding.
@@ I Introduction
  • A 12-25x reduction in the number of iterations needed to train the neural label field and practical runtime improvement of 14-24$\times$ over prior art, benchmarked on a single NVIDIA RTX 3090.
@@ II Background
2D segmentation: The prevalence of vision transformer architecture and the increasing scale of image datasets have resulted in a series of state-of-the-art image segmentation models. Panoptic and Contrastive Lifting both lift panoptic segmentation masks produced by Mask2Former [1] to 3D by learning a neural field. Towards open-set segmentation, segment anything (SAM) [2] achieves unprecedented performance by training on a billion masks over 11 million images. HQ-SAM [9] improves upon SAM for fine-grained masks. FastSAM [8] distills SAM into a CNN architecture and achieves similar performance while being orders of magnitude faster. In this work, we use GroundedSAM [10, 11], which combines SAM with Recognize Anything (RAM) [10] and GroundingDino [11] to produce object-level, instead of part-level segmentation masks.
Neural fields for 3D instance segmentation: NeRFs are implicit scene representations that can accurately encode complex geometry, semantics, and other modalities and resolve viewpoint inconsistent supervision [12]. Panoptic lifting [5] constructs semantics and instances branches on an efficient variant of NeRF, TensoRF [13], utilizing a Hungarian matching loss function to assign learned instance masks to surrogate object IDs given reference view-inconsistent masks. This scales poorly with an increasing number of objects (owing to the cubic complexity of Hungarian matching). Contrastive lifting [6] addresses this by instead employing contrastive learning on scene features, with positive and negative relations determined by whether or not they project onto the same mask. In addition, contrastive lifting requires a slow-fast clustering-based loss for stable training, leading to faster performance than panoptic lifting but requires multiple stages of training, leading to slow convergence. Concurrently to us, Instance-NeRF [14] directly learns a label field via supervision, but their instance labels are derived from the 3D object detector NeRF-RPN [15], which cannot detect small objects. On the contrary, our approach is based on lifting SAM, which allows us to segment objects at SAM’s granularity.
Structure from Motion: During mask association in InstanceMap, we take inspiration from scalable 3D reconstruction pipelines such as Hierarchical Localization (hLoc) [16]. Specifically, we use visual descriptors produced by NetVLAD for matching image viewpoints as a first step, then perform keypoint extraction and matching using LoFTR [17] as a preliminary for mask association on the matched image viewpoints.
@@ IV-C Implementation Details
For comparison with Panoptic and Contrastive Lifting, we follow their evaluation framework and map the Mask2Former outputs and ground truth semantic labels to 21 reduced ScanNet classes, considering only the "thing" subset of those classes. We train Panoptic and Contrastive Lifting using the code provided by [6]. We use the default parameterizations as their Replica experiment, which scales the number of iterations (128000) with the number of images. We build 3DIML on top of the popular framework Nerfstudio [18] with tinycudann enabled [19]. Training using the pseudolabels produced by InstanceMap allows us to dramatically reduce the number of training iterations to 5000 iterations for Replica-vMap (both Mask2Former and GroundedSAM) and 10000 iterations for ScanNet. For our panoptic experiments, we use $\left(\tau_{\text{global}},\tau_{\text{local}},\tau_{\text{community}},\tau_{\text{area}},\tau_{\text{merge}}\right)=$ (0.25, 0.8, 2, 0.15, 0.75). For GroundedSAM, we use (0.25, 0.75, 3, 0.3, 0.6) for Replica-vMap and (0.4, 0.75, 3, 0.25, 0.75) for ScanNet.